# From quantum enhanced to quantum inspired Monte Carlo

### v1

@@ Abstract
We perform a comprehensive analysis of the quantum-enhanced Monte Carlo method [Nature, 619, 282–287 (2023)], aimed at identifying the optimal working point of the algorithm. We observe an optimal mixing Hamiltonian strength and analyze the scaling of the total evolution time with the size of the system. We also explore extensions of the circuit, including the use of time-dependent Hamiltonians and reverse digitized annealing. Additionally, we propose that classical, approximate quantum simulators can be used for the proposal step instead of the original real-hardware implementation. We observe that tensor-network simulators, even with unconverged settings, can maintain a scaling advantage over standard classical samplers. This may extend the utility of quantum enhanced Monte Carlo as a quantum-inspired algorithm, even before the deployment of large-scale quantum hardware.
@@ I Introduction
Very recently, with the advent of quantum computation [5], efforts have been made to devise quantum algorithms that could speed up the sampling tasks of classical lattice models [6]. Quantum walks were proposed decades ago [7], but they require a long coherent evolution, and it is not clear whether the quadratic speed-up they offer is enough to provide a real practical advantage [8, 9]. More recently, it has been proposed that wavefunction collapses could be a powerful computational resource for uncorrelated configurations sampling in physical models [10]. This general idea has been later formalized in the quantum-enhanced Markov Chain Monte Carlo algorithm (QeMC), which features a hybrid strategy [11]: the proposal step of the Markov chain is performed using quantum hardware, while the acceptance step, which requires evaluating the cost-function difference between the current and proposed configuration, is done classically. The quantum update is realized using a Hamiltonian evolution circuit, implemented through the Trotter algorithm. The main advantage of this method is that coherent evolution is not required throughout the entire Markov chain, making it a plausible candidate for near-term quantum advantage. Moreover, the method displays an empirical, super-quadratic advantage over the best local update Metropolis scheme for average-case disordered instances, although this evidence is currently limited to lattice sizes of up to about 20. The algorithm has been demonstrated on real quantum hardware, displaying an interesting resilience to hardware noise. However, the evidence for quantum advantage is empirical, and a clear understanding of why the method works is still lacking. This is, however, common in the field of Monte Carlo, where the efficiency of sampling schemes is ultimately proven numerically.
@@
Multiple, recent studies have been dedicated to better understanding the algorithm in exactly solvable, specific models [12] or parameter’s limits [13], proposing alternative circuit implementations [14], extensions to continuous models [15], or further exploring parallelization possibilities [16].
@@
In this manuscript, we focus on two major points. The first concerns the sensitivity of the quantum speed-up with respect to the QeMC algorithm’s parameters, namely the strength of the mixing Hamiltonian and the total evolution time (see Sect. II). This analysis is important to determine whether the algorithm requires excessive instance-dependent fine-tuning, which would make it impractical, and whether there is a hidden relationship between the simulation time and the system size. Clearly, if the depth of the circuit needs to grow exponentially with the system size, this would diminish the quantum advantage.
@@
The second, conceptually novel, point is to explore whether a quantum-inspired version of the algorithm, running on classical machines, can be also effective. While classical methods cannot exactly simulate quantum dynamics, QeMC uses the Hamiltonian simulation circuit only as an update proposal generator. It remains unclear whether algorithmic errors can significantly impact the overall quantum speed-up. For instance, it is known that running simulated quantum annealing [17, 18] (via path integral Monte Carlo simulations) in an unconverged, unphysical settings can actually be beneficial if this method is used for optimization. In such cases, large imaginary-time Trotter steps [19], or open imaginary-time boundary conditions [20, 21] improve the probability of escaping local minima. Therefore, it is not ruled out that using the Hamiltonian circuit with a larger Trotter step, or even more radically, employing a classical approximation of quantum dynamics within QeMC, could preserve – or possibly boost – its efficiency.
@@
The paper is organized as follows. In Sect. II we introduce the original quantum-enhance Monte Carlo algorithm and state the requirements fo practical quantum advantage. In Sect. III we identify the optimal working point of the algorithm. In Sect. IV we test different circuits for the proposal step. In Sect. V we introduce the quantum inspired version of the technique. We draw our conclusions and outlook in Sect. VI.
@@ II.1 Spin glasses
Let $H_{c}(s)$ be a classical Ising model Hamiltonian and $s\in\{\pm 1\}^{n}$ be a classical state in the configuration space of a system of $n$ spin variables. An Ising model is defined by coefficients $\{J_{ij}\}$ and $\{h_{i}\}$ called couplings and fields, respectively. A configuration $s$ has energy
@@
In this paper, we will consider a fully connected graph such that all $J_{ij}$ are nonzero. We focus on the fully connected model, as it is more challenging. The corresponding Boltzman distribution $\pi(s)$ at temperature $T=1/\beta$ is given by
@@
where $\mathcal{Z}=\sum_{s}{e^{-\beta H_{c}(s)}}$ is the normalization constant. Usually, computing the Boltzmann distribution explicitly is computationally intractable due to an exponentially large summation in the normalization constant. Ising models with random ${J_{ij}}$ and ${h_{i}}$, also known as spin glasses, typically exhibit rugged energy landscapes with many local minima [22, 23, 24, 25]. While this type of Hamiltonian may seem artificial, the model is ubiquitous across many fields of science and engineering, from materials science, and optimization, to biological networks [26, 27]. Historically, disordered spin glasses have been considered an ideal testbed for developing and testing novel classical [28, 17, 29, 30, 31] and quantum algorithms and hardware [32, 33, 34, 35, 36, 11].
@@ II.2 Markov chain Monte Carlo
Markov chain Monte Carlo (MCMC) is the most widely used technique for sampling from intractable Boltzmann distributions [37, 38]. MCMC avoids explicit computation of $\pi(s)$ through an efficient two-step process. Initially, starting from a spin configuration $s_{0}$, a new configuration $s$ is proposed with a probability $Q(s|s_{0})$. Subsequently, the proposed configuration is accepted with probability $A(s|s_{0})$. These two steps form a Markov chain, whereby the transition from $s_{0}$ to $s$ occurs with a probability
@@
$A(s|s_{0})=\min\left(1,\frac{\pi(s)Q(s_{0}|s)}{\pi(s_{0})Q(s|s_{0})}\right)$ (4)
@@
The performance of MCMC is evaluated by its convergence rate to the stationary distribution $\pi$. Fast convergence is essential for practical applications. The convergence rate is defined by the mixing time $t_{mix}$, which is the number of steps required for the chain to become $\epsilon$-close (in total variation distance) to the stationary distribution. The distribution at any given time step is determined by the repeated application of the transition matrix $P$. The stationary distribution corresponds to the eigenvalue $\lambda_{1}=1$ of the transition matrix, and the spectral gap determines the convergence rate [39]:
$\delta=1-\max_{i\neq 1}|\lambda_{i}|$ (5)
where $1=\lambda_{1}\geq\lambda_{2}\geq\lambda_{3}\geq\ldots\geq\lambda_{d}$ are the eigenvalues of $P$. The mixing time can be bounded by the spectral gap as follows:
@@
It is important to note that although the spectral gap is an unambiguous metric to compare different MCMC methods, calculating $\delta$ requires diagonalizing the transition matrix. This process is much more computationally demanding than just running MCMC and is typically only feasible for small systems.
@@ II.3 Quantum-enhanced Markov chain Monte Carlo
The acceptance step remains identical to that in the classical MCMC. However, significant simplification is achieved by selecting a symmetric operator $U$. When $U=U^{T}$, the Metropolis-Hastings acceptance rule reduces to:
@@
The original work considers the evolution under a time-independent quantum Hamiltonian of the spin glass in a transverse field, often referred to as the quantum Sherrington-Kirkpatrick model or quantum spin glass [40, 41, 42]. The Hamiltonian is given by:
@@
where $H_{\textrm{mix}}$ is a mixing Hamiltonian with non-zero off-diagonal elements to generate transitions. Specifically, $H_{\textrm{mix}}$ is defined as:
@@
The scale factor $\alpha=\frac{||H_{\textrm{mix}}||_{F}}{||H_{c}||_{F}}$ establishes a common energy scale between the two components of the Hamiltonian, where ||..$||..||_{F}$ is the Frobenius norm. Finally, the parameter $\gamma$ determines the relative weight of the mixing Hamiltonian. Consequently, the time evolution operator is given by
@@
where $t$ represents the total evolution time. Since both terms in the Hamiltonian in Eq. 9 are symmetric, the time evolution operator $U$ is also symmetric. Practically, $U$ can be approximated using a second-order Suzuki-Trotter expansion [11].
In the original Ref. [11], the parameters, $\gamma$ and $t$, are selected uniformly at random for each proposal step. This heuristic approach eliminates the need for optimization and is based on the idea that even if $\gamma$ and $t$ are occasionally chosen suboptimally, the algorithm will still perform adequately as long as optimal values are selected frequently enough.
Finally, we note that the choice of the mixing Hamiltonian is arbitrary, as long as $U$ remains symmetric and the term allows for ergodic exploration of the configurational space. This holds true for the simple one-body transverse field. In the earlier Ref. [10], the idea was introduced in the continuum, where an unambiguous choice for the mixing (or kinetic) term is possible: the Fokker-Planck Hamiltonian, which fully determines the equilibrium and kinetic properties of a system. In principle, as we will see in Sect. IV.2, one could go beyond this structure, and employ unitaries which are not strictly defined by Hamiltonian simulation [14].
Figure 1: (a) Grid search over $\gamma$ and $t$ for the spectral gap $\delta$ for $n=9$ system. Throughout this study, each data point is an average over 100 model instances and $T=1$. Warmer colors indicate higher $\delta$, i.e., faster MCMC convergence. (b) Spectral gap $\delta_{avg}$ averaged over $t\in[12,30]$ as a function of $\gamma$. The dashed red lines represent systems $n=4$ to $n=8$ and $n=9$ is represented by a solid line. (c) Optimal gamma values $\gamma_{\textrm{opt}}$ plotted as a function of system size. Error bars represent the standard deviation.
@@ II.4 Thresholds for practical quantum advantage
The main result of Ref. [11] is that QeMC is able to achieve a scaling advantage for the mixing time compared to all possible classical, local and cluster Monte Carlo update schemes. In particular, the dependence of the gap $\delta$, Eq. 5, on the system size, $n$ is analyzed. The gap closes exponentially (meaning that the mixing time increases exponentially), as $\delta\propto 2^{-kn}$, with $k>0$, for all schemes considered. However the QeMC exponent is $k_{q}=0.264(4)$, while the classical sampler shows $k_{c}=0.94(4)$, resulting in a polynomial speed-up factor of $\alpha=k_{c}/k_{q}=3.6(1)$ (cfn. also Fig 3). This means that QeMC features smaller autocorrelation times, allowing it to achieve a target sampling quality with fewer samples compared to classical MCMC, which has a longer mixing time.
[…]
@@
Therefore, the ratio of the quantum and classical runtimes per iteration, $t_{q}/t_{c}$, is logarithmically suppressed compared to the inverse difference between the scaling exponents.
We attempt a rough, order-of-magnitude estimate of the conditions that must be met for a practical advantage on a concrete case study. We follow the discussion in Lemieux et. al. [8], where quantum walks were studied. First, it is necessary to identify the classical state of the art and a problem size that is challenging for current methods. In Ref. [8], the Janus FPGA-based special-purpose machine is selected as a benchmark. This hardware was built to simulate a $80\times 80\times 80$ cubic spin glass with local couplings and can perform about one spin flip per picosecond. The numerical experiments involve a total of $10^{18}$ local updates (spin flip) and run for about one month [44].
These parameters establish a threshold for quantum advantage under different scenarios. Assuming that QeMC has a true scaling advantage of order $\alpha$, this means it could achieve the same sampling quality, i.e., accuracy on computing thermodynamic averages, with $10^{18/\alpha}$ samples. For the sake of concreteness, let’s assume values of $\alpha=4,3,2$. For $\alpha=4$, i.e. quartic speedup, we find that the crossover for quantum advantage occurs at outputting $10^{4.5}$ samples in one month, which means performing about one quantum proposal step per minute. Conversely, if $\alpha=3$, it would require performing $10^{6}$ QeMC steps in one month, meaning one step every 2.5 seconds. The observed speed-up in Ref. [11], 3.6(1), stands in between these values. Finally, if the speed-up is only quadratic, the possibility of practical quantum advantage diminishes, as it would require performing one QeMC step every 2 milliseconds.
While these constraints seem within reach, the quantum circuit in this case is defined over about half a million qubits. In this example, the system’s size is very large, but the couplings are local. Fully connected models become hard at much lower sizes [25]
Another critical consideration is the scaling of the Hamiltonian simulation time, $t$ in Eq. 11, and consequently the circuit depth, with the system size, $n$. Obviously, a scaling dependency $t\propto 2^{n\beta}$ will decrease the order of quantum speedup from $\alpha=k_{c}/k_{q}$ to $\alpha=k_{c}/(k_{q}+\beta)$. However, in Sect.III.2 we show that the total evolution time likely scales subexponentially, and thus does not affect asymptotic quantum speedup.

### v2

# thanks: equal contributionthanks: equal contribution

@@ Abstract
We perform a comprehensive analysis of the quantum-enhanced Monte Carlo method [Nature, 619, 282 (2023)], aimed at identifying the optimal working point of the algorithm. We observe an optimal mixing Hamiltonian strength and analyze the scaling of the total evolution time with the size of the system. We also explore extensions of the circuit, including the use of time-dependent Hamiltonians and reverse digitized annealing. Additionally, we propose that classical, approximate quantum simulators can be used for the proposal step instead of the original real-hardware implementation. We observe that tensor-network simulators, even with unconverged settings, can maintain a scaling advantage over standard classical samplers. This may extend the utility of quantum enhanced Monte Carlo as a quantum-inspired algorithm, even before the deployment of large-scale quantum hardware.
@@ I Introduction
Very recently, with the advent of quantum computation [5], efforts have been made to devise quantum algorithms that could speed up the sampling tasks of classical lattice models [6]. Quantum walks were proposed decades ago [7], but they require a long coherent evolution, and it is not clear whether the quadratic speed-up they offer is enough to provide a real practical advantage [8, 9]. More recently, it has been proposed that wavefunction collapses could be a powerful computational resource for uncorrelated configurations sampling in physical models [10]. This general idea was later formalized in the quantum-enhanced Markov chain Monte Carlo (QEMC) algorithm, which features a hybrid strategy [11]: The proposal step of the Markov chain is performed using quantum hardware, while the acceptance step, which requires evaluating the cost-function difference between the current and proposed configurations, is done classically. The quantum update is realized using a Hamiltonian evolution circuit, implemented through the Trotter algorithm. The main advantage of this method is that coherent evolution is not required throughout the entire Markov chain, making it a plausible candidate for near-term quantum advantage. Moreover, the method displays an empirical superquadratic advantage over the best local update Metropolis scheme for average-case disordered instances, although this evidence is currently limited to lattice sizes of up to about 20. The algorithm has been demonstrated on real quantum hardware, displaying an interesting resilience to hardware noise. However, the evidence for quantum advantage is empirical, and a clear understanding of why the method works is still lacking. This is, however, common in the field of Monte Carlo, where the efficiency of sampling schemes is ultimately proven numerically.
@@
Multiple recent studies have been dedicated to better understanding the algorithm in exactly solvable, specific models [12] or the parameter’s limits [13], proposing alternative circuit implementations [14], extensions to continuous models [15], or further exploring parallelization possibilities [16].
@@
In this paper we focus on two major points. The first concerns the sensitivity of the quantum speed-up with respect to the QEMC algorithm’s parameters, namely the strength of the mixing Hamiltonian and the total evolution time (see Sec. II). This numerical analysis is important to determine whether the algorithm requires excessive instance-dependent fine-tuning, which would make it impractical, and whether there is a hidden relationship between the simulation time and the system size. Clearly, if the depth of the circuit needs to grow exponentially with the system size, this would diminish the quantum advantage.
@@
The second, conceptually novel, point is to explore whether a quantum-inspired version of the algorithm, running on classical machines, can also be effective. While classical methods cannot exactly simulate quantum dynamics, QEMC uses the Hamiltonian simulation circuit only as an update proposal generator. It remains unclear whether algorithmic errors can significantly impact the overall quantum speedup. For instance, it is known that running simulated quantum annealing [17, 18] (via path-integral Monte Carlo simulations) in an unconverged, unphysical setting can actually be beneficial if this method is used for optimization. In such cases, large imaginary-time Trotter steps [19] or open imaginary-time boundary conditions [20, 21] improve the probability of escaping local minima. Therefore, it is not ruled out that using the Hamiltonian circuit with a larger Trotter step, or even more radically, employing a classical approximation of quantum dynamics within QEMC, could preserve, or possibly boost, its efficiency.
@@
The paper is organized as follows. In Sec. II we introduce the original quantum-enhanced Monte Carlo algorithm and state the requirements fo practical quantum advantage. In Sec. III we identify the optimal working point of the algorithm. In Sec. IV we test different circuits for the proposal step. In Sec. V we introduce the quantum-inspired version of the technique. In Sec. VI we summarize the work, discuss our conclusions, and present an outlook for future work.
@@ II.1 Spin glasses
Let $H_{c}(s)$ be a classical Ising model Hamiltonian and $s\in\{\pm 1\}^{n}$ be a classical state in the configuration space of a system of $n$ spin variables. An Ising model is defined by coefficients $\{J_{ij}\}$ and $\{h_{i}\}$, called couplings and fields, respectively. A configuration $s$ has energy
@@
In this paper, we consider a fully connected graph such that all $J_{ij}$ are nonzero. We focus on the fully connected model, as it is more challenging. The corresponding Boltzmann distribution $\pi(s)$ at temperature $T=1/\beta$ is given by
@@
where $\mathcal{Z}=\sum_{s}{e^{-\beta H_{c}(s)}}$ is the normalization constant. Usually, computing the Boltzmann distribution explicitly is computationally intractable due to an exponentially large summation in the normalization constant. Ising models with random ${J_{ij}}$ and ${h_{i}}$, also known as spin glasses, typically exhibit rugged energy landscapes with many local minima [22, 23, 24, 25]. While this type of Hamiltonian may seem artificial, the model is ubiquitous across many fields of science and engineering, from materials science and optimization to biological networks [26, 27]. Historically, disordered spin glasses have been considered an ideal testbed for developing and testing novel classical [3, 17, 28, 29, 30] and quantum algorithms and hardware [31, 32, 33, 34, 35, 11].
@@ II.2 Markov chain Monte Carlo
Markov chain Monte Carlo (MCMC) is the most widely used technique for sampling from intractable Boltzmann distributions [36, 37]. It avoids explicit computation of $\pi(s)$ through an efficient two-step process. Initially, starting from a spin configuration $s_{0}$, a new configuration $s$ is proposed with a probability $Q(s|s_{0})$. Subsequently, the proposed configuration is accepted with probability $A(s|s_{0})$. These two steps form a Markov chain, whereby the transition from $s_{0}$ to $s$ occurs with a probability
@@
$A(s|s_{0})=\min\left(1,\frac{\pi(s)Q(s_{0}|s)}{\pi(s_{0})Q(s|s_{0})}.\right)$ (4)
@@
The performance of MCMC is evaluated by its convergence rate to the stationary distribution $\pi$. Fast convergence is essential for practical applications. The convergence rate is defined by the mixing time $t_{\text{mix}}$, which is the number of steps required for the chain to become $\epsilon$-close (in total variation distance) to the stationary distribution. The distribution at any given time step is determined by the repeated application of the transition matrix $P$. The stationary distribution corresponds to the eigenvalue $\lambda_{1}=1$ of the transition matrix, and the spectral gap determines the convergence rate [38]
$\delta=1-\max_{i\neq 1}|\lambda_{i}|,$ (5)
where $1=\lambda_{1}\geq\lambda_{2}\geq\lambda_{3}\geq\cdots\geq\lambda_{d}$ are the eigenvalues of $P$. The mixing time can be bounded by the spectral gap as follows:
@@
It is important to note that although the spectral gap is an unambiguous metric to compare different MCMC methods, calculating $\delta$ requires diagonalizing the Hamiltonian $H$ to compute the proposal matrix $Q$ and the dense transition matrix $P$. The computational complexity associated with diagonalizing $P$ is $O\left(2^{3n}\right)$. Therefore, our analysis is limited to small systems $n<10$. These sizes are consistent with the ones originally investigated in Ref. [11]. Note that storing the full transfer matrix $P$ is obviously not necessary in practical Monte Carlo simulations. However, this remains the most accurate method for computing mixing times. To maintain full consistency with the original work in [11], we also adopt this strategy. In other words, in this numerical setup, we choose to calculate the mixing times accurately by constructing and diagonalizing an exponentially large, dense matrix. However, in practical Metropolis simulations, whether quantum or quantum-inspired, this will not be necessary.
@@ II.3 Quantum-enhanced Markov chain Monte Carlo
The acceptance step remains identical to that in the classical MCMC. However, significant simplification is achieved by selecting a symmetric operator $U$. When $U=U^{T}$, the Metropolis-Hastings acceptance rule reduces to
@@
The original work considers the evolution under a time-independent quantum Hamiltonian of the spin glass in a transverse field, often referred to as the quantum Sherrington-Kirkpatrick model or quantum spin glass [39, 40, 41]. The Hamiltonian is given by
@@
where $H_{\textrm{mix}}$ is a mixing Hamiltonian with non-zero off-diagonal elements to generate transitions. Specifically, $H_{\textrm{mix}}$ is defined as
@@
The scale factor $\alpha=\frac{||H_{\textrm{mix}}||_{F}}{||H_{c}||_{F}}$ establishes a common energy scale between the two components of the Hamiltonian, where $\|\cdots\|_{F}$ is the Frobenius norm. Finally, the parameter $\gamma$ determines the relative weight of the mixing Hamiltonian. Consequently, the time-evolution operator is given by
@@
where $t$ represents the total evolution time. Since both terms in the Hamiltonian in Eq. 9 are symmetric, the time-evolution operator $U$ is also symmetric. Practically, $U$ can be approximated using a second-order Suzuki-Trotter expansion [11].
In Ref. [11] the parameters $\gamma$ and $t$ are selected uniformly at random for each proposal step. This heuristic approach eliminates the need for optimization and is based on the idea that even if $\gamma$ and $t$ are occasionally chosen suboptimally, the algorithm will still perform adequately as long as optimal values are selected frequently enough.
Finally, we note that the choice of the mixing Hamiltonian is arbitrary, as long as $U$ remains symmetric and the term allows for ergodic exploration of the configurational space. This holds true for the simple one-body transverse field. In Ref. [10] the idea is introduced in the continuum, where an unambiguous choice for the mixing (or kinetic) term is possible: the Fokker-Planck Hamiltonian, which fully determines the equilibrium and kinetic properties of a system. In principle, as we will see in Sec. IV.2, one could go beyond this structure and employ unitaries which are not strictly defined by Hamiltonian simulation [14].
Figure 1: (a) Grid search over $\gamma$ and $t$ for the spectral gap $\delta$ for the $n=9$ system. Throughout this study, each data point is an average over 100 model instances and $T=1$. Warmer colors indicate higher $\delta$, i.e., faster MCMC convergence. (b) Spectral gap $\delta_{\text{avg}}$ averaged over $t\in[12,30]$ as a function of $\gamma$. The red dashed lines represent systems $n=4-8$, and the red solid line represents $n=9$. (c) Optimal gamma values $\gamma_{\textrm{opt}}$ plotted as a function of system size. Error bars represent the standard deviation.
@@ II.4 Thresholds for practical quantum advantage
The main result of Ref. [11] is that QEMC is able to achieve a scaling advantage for the mixing time compared to all possible classical, local, and cluster Monte Carlo update schemes. In particular, the dependence of the gap $\delta$ [Eq. (5)] on the system size $n$ is analyzed. The gap closes exponentially (meaning that the mixing time increases exponentially), as $\delta\propto 2^{-kn}$, with $k>0$, for all schemes considered. However, the QEMC exponent is $k_{q}=0.264(4)$, while the classical sampler shows $k_{c}=0.94(4)$, resulting in a polynomial speed-up factor of $\alpha=k_{c}/k_{q}=3.6(1)$ (see also Fig. 3). This means that QEMC features smaller autocorrelation times, allowing it to achieve a target sampling quality with fewer samples compared to classical MCMC, which has a longer mixing time.
Unfortunately, the theoretical scaling advantage can still be overshadowed by the large gate-time prefactor, even for reasonably large system sizes [9]. Note that the time required to execute a quantum gate is several orders of magnitude slower than that of a classical one. Accepted estimates for the gate clock of future fault-tolerant hardware are on the order of 10 kHz [42]. This implies that the quantum algorithm starts at a significant disadvantage compared to a classical machine, which is not compensated by more favorable scaling until a certain crossover problem size is reached. Current noisy machines operate at higher frequencies but lack error correction. More precisely, the classical and quantum runtimes are $\mathcal{T}_{c}(n)\propto t_{c}(n)~{}2^{k_{c}n}$ and $\mathcal{T}_{q}(n)\propto t_{q}(n)~{}2^{k_{q}n}$, respectively, where $t_{c/q}(n)$ are the runtimes to perform one single MCMC step in both schemes. This includes not only the algorithmic scaling with the system size but also the hardware gate-time prefactor. The crossover size at which the quantum algorithm outperforms the classical one is at
@@
Therefore, the ratio of the quantum and classical runtimes per iteration $t_{q}/t_{c}$ is logarithmically suppressed compared to the inverse difference between the scaling exponents.
We attempt a rough, order-of-magnitude estimate of the conditions that must be met for a practical advantage on a concrete case study. We follow the discussion in Ref. [8], where quantum walks are studied. First, it is necessary to identify the classical state of the art and a problem size that is challenging for current methods. In Ref. [8] the Janus field-programmable gate array–based special-purpose machine is selected as a benchmark. This hardware was built to simulate a $80\times 80\times 80$ cubic spin glass with local couplings and can perform about one spin flip per picosecond. The numerical experiments involve a total of $10^{18}$ local updates (spin flip) and run for about one month [43].
These parameters establish a threshold for quantum advantage under different scenarios. Assuming that QEMC has a true scaling advantage of order $\alpha$, this means it could achieve the same sampling quality, i.e., accuracy on computing thermodynamic averages, with $10^{18/\alpha}$ samples. For the sake of concreteness, let us assume values of $\alpha=4,3,2$. For $\alpha=4$, i.e. quartic speedup, we find that the crossover for quantum advantage occurs at outputting $10^{4.5}$ samples in one month, which means performing about one quantum proposal step per minute. Conversely, if $\alpha=3$, it would require performing $10^{6}$ QEMC steps in one month, meaning one step every 2.5 s. The observed speed up in Ref. [11], 3.6(1), stands in between these values. Finally, if the speed up is only quadratic, the possibility of practical quantum advantage diminishes, as it would require performing one QEMC step every 2 ms.
While these constraints seem within reach, the quantum circuit in this case is defined over about 500,000 qubits. In this example, the system’s size is very large, but the couplings are local. Fully connected models become hard at much lower sizes [25]
Another critical consideration is the scaling of the Hamiltonian simulation time $t$ in Eq. (11), and consequently the circuit depth, with the system size, $n$. Obviously, a scaling dependence $t\propto 2^{n\beta}$ will decrease the order of quantum speedup from $\alpha=k_{c}/k_{q}$ to $\alpha=k_{c}/(k_{q}+\beta)$. However, in Sec.III.2 we show that the total evolution time likely scales subexponentially and thus does not affect asymptotic quantum speedup.